# Assessing the Feasibility of Internet-Sourced Video for Automatic Cattle Lameness Detection

### v1

@@ Abstract
Cattle lameness is often caused by hoof injuries or interdigital dermatitis, leads to pain and significantly impacts essential physiological activities such as walking, feeding, and drinking. This study presents a deep learning-based model for detecting cattle lameness, sickness, or gait abnormalities using publicly available video data. The dataset consists of 50 unique videos from 40 individual cattle, recorded from various angles in both indoor and outdoor environments. Half of the dataset represents naturally walking (normal/non-lame) cattle, while the other half consists of cattle exhibiting gait abnormalities (lame). To enhance model robustness and generalizability, data augmentation was applied to the training data. The pre-processed videos were then classified using two deep learning models: ConvLSTM2D and 3D CNN. A comparative analysis of the results demonstrates strong classification performance. Specifically, the 3D CNN model achieved a video-level classification accuracy of 90%, with precision, recall, and f1-score of 90.9%, 90.9%, and 90.91% respectively. The ConvLSTM2D model exhibited a slightly lower accuracy of 85%. This study highlights the effectiveness of directly applying classification models to learn spatiotemporal features from video data, offering an alternative to traditional multi-stage approaches that typically involve object detection, pose estimation, and feature extraction. Besides, the findings demonstrate that the proposed deep learning models, particularly the 3D CNN, effectively classify and detect lameness in cattle while simplifying the processing pipeline.
@@
Cattle, Lameness detection, Deep-learning, Computer vision techniques, Image processing
@@ I Introduction
Cattle farms significantly contribute to global food security by providing meat and dairy products, with continuously increasing demand. While large-scale farming is now common, it presents operational challenges and necessitates continuous monitoring of animal health and welfare, as this directly impacts farm productivity [1, 2]. However, one of the leading health issues in cattle is lameness, a common and painful disorder frequently caused by gait-related injuries. Research indicates that approximately one in four cattle is affected by lameness at any given time, which reduces farm productivity and plays a substantial role in increasing meat and dairy prices [3, 5, 4]. A common practice about identifying lame cattle is herd locomotion scoring which requires visual observation of each animal individually. This process is cost and time consuming since they tend to increase human commissioning [6, 7]. Therefore, it is highly necessary to deploy advanced automatic technologies for identification of such health difficulties.
In recent years, AI through ML and data analysis have come up with smart solutions to serve large-scale cattle farming in different ways, e.g., by early detection of diseases and infections, by improving the production characteristics, etc [8]. AI uses data collected from livestock farms using various advanced tools including cameras, drones, robotics, and relevant IoT devices [9]. Although manual video cameras have been commonly used for a long time to surveillance the activities of animals from farmers’ desks, nowadays, computer vision and pattern recognition techniques have added new opportunities using the video data to automatically monitor animals’ health and welfare 24/7. Whether using a manual or computer vision (CV) approach, lameness is detected by observing behavioral changes in cattle—how they react and respond during regular activities in their living environment. Advanced deep learning-oriented AI models analyze image or video data from different perspectives, depending on the animal’s behavior (e.g., eating, walking, and general movement) [10]. Abnormal behavior is often readily apparent when cattle walk, as lameness is closely associated with painful gait and foot disease. The automation primarily focuses on detecting deviations between normal and abnormal behavior, classifying them to make decisions and further analysis.
@@ II Materials and Methods
TABLE I: Basic information of the dataset used in this study with the most relevant datasets in the literature
Study Dataset Size Test Environment Focused Activity
Train Frame Test Frame
@@
5 600 120 The corridor after the milking parlor Intensive farming
@@
This study 3000 1000 Indoor and outdoor environment Semi-Intensive farming
II-A Data
II-A1 Data collection
The data collection process for this study was quite different from that of previous studies. We used an online platform to collect freely accessible cattle walking videos, specifically from YouTube. Publicly available data are commonly used in many research fields; for example, one of our previous articles [15] demonstrated that YouTube videos are frequently used in deepfake video detection research. In order to create an inclusive classification model, we created a diverse dataset. Therefore, the following features were incorporated with the video: there were different cattle in different environments, including intensive, semi-intensive, and extensive cattle farming; the cattle walked in different directions, including left to right and right to left; and the videos were captured from various views, including front, side, and back. For this study, a total of 50 videos featuring 40 cattle were considered, with an average duration of five seconds per video. We ensured an equal class distribution of the data, meaning that 25 lame and 25 non-lame cattle videos were included.
@@ II-A1 Data collection
Table I provides an overview of the dataset used in this research alongside datasets from previous cattle lameness detection studies. The table compares these studies based on the number of training and testing frames, the test environment, and the focused activity. As shown in the last row, this study employs a comparatively larger dataset, consisting of 3,000 training and 1,000 testing frames. Previous studies focused on intensive farming environments, where cattle were primarily managed in indoor grazing conditions and data was collected from specific, limited locations such as within a barn, a passing alley, a narrow path, and corridors related to milking parlors. In contrast, this study focuses on a semi-intensive farming setting and incorporates both indoor and outdoor environments. This approach captures a wider range of cattle movement behaviors, lighting conditions, and surface variations, potentially leading to a more generalizable model for cattle lameness detection.
II-A2 Data pre-processing
@@
Preprocessing in this investigation was conducted in several distinct steps, including frame extraction and data augmentation. Initially, frames were extracted at the standard rate of 30 frames per second. However, due to the high volume of frames, only 50 images were randomly selected from each training video, resulting in a total of 1500 images. These images were then cropped to a 608×608 pixel ratio. However, data augmentation is a widely recognized method for addressing limited data issues and helps to create robust and sustainable classification models. Numerous data augmentation techniques are available, as demonstrated by Jia et al. [3], who employed nine different methods, such as Mixup, Cutout, CutMix, etc. This study considered only image flipping as an augmentation technique, applied solely to the training data. All 30 training videos (1500 frames) were virtually flipped, generating a total of 3000 images.
II-B Model Construction
Algorithm 1 Pseudo-code of the proposed classification model
[…]
2:Trained 3DCNN OR ConvLSTM2D model; Performance metrics
3:1. Data Preparation:
4:for $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{test}}$ do
5:     $F_{i}\leftarrow\text{Pad/Truncate}(\text{ResizeFrames}(\text{ExtractFrames}(V_{i},\text{max\_frames}=50),S_{\text{target}}),N_{\text{frames}})$
6:     $(X,y)\leftarrow(X\cup\{F_{i}\},y\cup\{l_{i}\})$ if $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}$ else $(X_{\text{test}},y_{\text{test}})$
7:end for
8:Augment $X_{\text{train}},y_{\text{train}}$ with horizontal flipping (probability = $p_{\text{aug}}$).
9:Normalize and encode labels for $X_{\text{train}},y_{\text{train}},X_{\text{test}},y_{\text{test}}$.
10:2. Model Training:
11:$M\leftarrow\text{Create 3DCNN OR ConvLSTM2D Model}(N_{\text{frames}},S_{\text{target}})$
[…]
13:3. Model Evaluation:
14:$\hat{y}\leftarrow M(X_{\text{test}})$
15:metrics, confusion matrix $\leftarrow\text{Evaluate}(y_{\text{test}},\hat{y})$
16:return $M$, metrics, confusion matrix
@@ II-B Model Construction
II-B1 3D CNN
CNNs are widely used deep learning models for image processing tasks. 3D CNNs are an extension of these and can handle 3D data containing spatiotemporal information, such as videos. They utilize 3D convolutional kernels, which enable them to focus on spatial features extracted from input data and distinguish themselves from traditional 2D CNNs. Because of their ability to learn both spatial and temporal features, this study employed a basic 3D CNN model to classify cattle videos, where movement patterns are assessed over time. The model extracts features by processing a sequence of video frames as input data through two convolutional layers (each followed by a max-pooling layer). The extracted features are then transformed into a one-dimensional vector by a flattening layer which is further used in two fully connected dense layers with ReLU activation. Two dropout layers are used to prevent overfitting, Finally, a softmax output layer provides the probability distribution over the defined classes.
II-B2 ConvLSTM2D
ConvLSTM2D is a deep learning architecture that integrates the strengths of spatial feature extraction and temporal dependency modeling through the combination of CNNs and LSTM networks, respectively. This makes it particularly effective for understanding patterns and movements in sequential data, such as videos. The ConvLSTM2D model employed in this study begins with a ConvLSTM2D layer. The model also incorporates two MaxPooling3D layers, one Conv3D layer, and dropout layers. Similar to the previous model, a dense layer with softmax activation is used to classify cattle into ‘normal’ and ‘lame’ categories.
II-C Model Evaluation
@@ III Results
(a) 3DCNN
(b) ConvLSTM2D
Figure 2: Confusion Matrix of both models under the same test data
TABLE II: Lameness classification results of proposed 3D CNN and ConvLSTM2D
@@
The models generate predictions for each individual frame within each video in the test data. It is assumed that if the model correctly predicts the majority of frames in a video, then the entire video is considered correctly classified, despite the presence of a few incorrect frame-level predictions. The confusion matrices for both models are presented in Figure 2. As shown in Figures 2a and 2b, a small number of misclassifications were observed, including 1 false positives and 1 false negative for 3D CNN, and 1 false positives and 2 false negatives for ConvLSTM2D, out of the total 20 samples (10 normal and 10 lame).
The 3D CNN model demonstrated notable performance in classifying the cattle data. Table II indicates an accuracy of 90%, along with strong average precision, recall, and F1-scores. On the other hand, ConvLSTM2D achieved comparatively lower performance, with accuracy, precision, recall, and F1-score values of 85%, 90%, 81.82%, and 85.71%, respectively. Overall, the confusion matrices and classification reports highlight the models’ effectiveness in classifying healthy and defective cattle.
@@ IV Discussion
TABLE III: Performance comparison of this study with existing cattle lameness classification models
@@
  • - No specific data was found
@@
Table III presents a performance comparison between eight prior studies and the proposed model. The table evaluates models based on accuracy, precision, recall, and f1-score. The final row of the table displays the results of this study, which employed a 3D CNN model, while the preceding rows summarize the findings from other relevant studies that utilized various machine learning and deep learning classifiers. The model used in this study achieved a notable accuracy of 90%. More significantly, it demonstrated the highest precision among the studies where precision values were available, reaching 90.9%. This indicates that the model exhibits a strong ability to minimize false positives, correctly identifying positive cases with high confidence. Additionally, the 3D CNN model also yielded a robust F1-score of 90.91%. When compared to other approaches, the 3D CNN model in this study demonstrated competitive performance, whereas study [10] employed C3D-ConvLSTM achieved a slightly higher accuracy of 90.32% but a lower precision of 90.62%.
@@ V Conclusions
This paper presents a cattle lameness video dataset and deep learning-based methods for effectively distinguishing between lame and non-lame cattle from video data. Compared to existing datasets used for cattle lameness detection, this study collected data entirely from online sources, providing more comprehensive information across diverse environments, scenes, and activities. Additionally, two deep learning models, 3D CNN and ConvLSTM2D, were used to classify normal and lame cattle. From the dataset, 3,000 images were extracted for training the models, while 1,000 images were used for validation. The experimental results demonstrated improved or at least comparable performance to existing cattle lameness classification models. In conclusion, this study serves as a reference for detecting cattle lameness or behavior without relying on traditional multi-step approaches.

### v2

@@ Abstract
Cattle lameness is often caused by hoof injuries or interdigital dermatitis, leads to pain and significantly impacts essential physiological activities such as walking, feeding, and drinking. This study presents a deep learning-based model for detecting cattle lameness, sickness, or gait abnormalities using publicly available video data. The dataset consists of 50 unique videos from 42 individual cattle, recorded from various angles in both indoor and outdoor environments. Half of the dataset represents naturally walking (normal/non-lame) cattle, while the other half consists of cattle exhibiting gait abnormalities (lame). To enhance model robustness and generalizability, data augmentation was applied to the training data. The pre-processed videos were then classified using two modified deep learning models: ConvLSTM2D and 3D CNN. A comparative analysis of the results demonstrates strong classification performance. Specifically, the 3D CNN model achieved a video-level classification accuracy of 90%, with precision, recall, and f1-score of 90.9%, 90.9%, and 90.91% respectively. The ConvLSTM2D model exhibited a slightly lower accuracy of 85%. This study highlights the effectiveness of directly applying classification models to learn spatiotemporal features from video data, offering an alternative to traditional multi-stage approaches that typically involve object detection, pose estimation, and feature extraction. Besides, the findings demonstrate that the proposed deep learning models, particularly the 3D CNN, effectively classify and detect lameness in cattle while simplifying the processing pipeline.
@@
Cattle, Lameness detection, Deep-learning, Computer vision techniques, Image processing.
@@ I Introduction
Cattle farms significantly contribute to global food security by providing meat and dairy products, with continuously increasing demand. While large-scale farming is now common, it presents operational challenges and necessitates continuous monitoring of animal health and welfare, as this directly impacts farm productivity [1, 2]. However, one of the leading health issues in cattle is lameness, a common and painful disorder frequently caused by gait-related injuries. Research indicates that approximately one in four cattle is affected by lameness at any given time, which reduces farm productivity and plays a substantial role in increasing meat and dairy prices [3, 5, 4]. A common practice about identifying lame cattle is herd locomotion scoring which requires visual observation of each animal individually [6, 7, 8, 9]. This process is cost and time consuming since they tend to increase human commissioning. Therefore, it is highly necessary to deploy advanced automatic technologies for identification of such health difficulties.
In recent years, AI through ML and data analysis have come up with smart solutions to serve large-scale cattle farming in different ways, e.g., by early detection of diseases and infections, by improving the production characteristics, etc [10]. AI uses data collected from livestock farms using various advanced tools including cameras, drones, robotics, and relevant IoT devices [14]. Although manual video cameras have been commonly used for a long time to surveillance the activities of animals from farmers’ desks [11, 12], nowadays, computer vision and pattern recognition techniques have added new opportunities using the video data to automatically monitor animals’ health and welfare 24/7 [13]. Whether using a manual or computer vision (CV) approach, lameness is detected by observing behavioral changes in cattle—how they react and respond during regular activities in their living environment. Advanced deep learning-oriented AI models analyze image or video data from different perspectives, depending on the animal’s behavior (e.g., eating, walking, and general movement) [15, 16]. Abnormal behavior is often readily apparent when cattle walk, as lameness is closely associated with painful gait and foot disease. The automation primarily focuses on detecting deviations between normal and abnormal behavior, classifying them to make decisions and further analysis.
@@
The remainder of this paper is organized as follows: Section II presents the materials and methods, including data collection, preprocessing, and model construction. Section III discusses the results obtained from the proposed models. Section IV discusses the findings and compares them with existing studies, along with limitations and future directions. Finally, Section V concludes the paper.
@@ II Materials and Methods
II-A Dataset
II-A1 Data collection
The data collection process for this study was quite different from that of previous studies. We used an online platform to collect freely accessible cattle walking videos, specifically from YouTube. Publicly available data are commonly used in many research fields; for example, one of our previous articles [21] demonstrated that YouTube videos are frequently used in deepfake video detection research. In order to create an inclusive classification model, we created a diverse dataset. Therefore, the following features were incorporated with the video: there were different cattle in different environments, including intensive, semi-intensive, and extensive cattle farming; the cattle walked in different directions, including left to right and right to left; and the videos were captured from various views, including front, side, and back.
II-A2 Basic Information
TABLE I: Summary of the dataset statistics.
Class Video Count Duration (s) Frame Count
Normal 25 134.46 4050
Lame 25 148.91 5900
Total 50 283.37 9950
To create the dataset, a total of 50 videos featuring 42 cattle were considered, with an average duration of 5.67 seconds per video. We ensured an equal class distribution of the data, meaning that 25 lame and 25 nomal cattle videos were included. Table I presents the dataset distribution, showing the number of videos, total duration in seconds, and the corresponding frame counts for the normal and lame classes. The normal class videos have a cumulative duration of 134.46 seconds and consist of 4050 frames, whereas the lame class videos have a total duration of 148.91 seconds and comprise 5900 frames. Overall, the dataset includes 50 videos with a combined duration of 283.37 seconds and a total of 9950 frames. Besides, Figure LABEL:fig:all_images displays example frames from cattle video sequences, with upper three images representing lame cattle and the other representing normal cattle. The original videos had varying resolutions due to collection from different online sources. To standardize the input data, all videos were resized to a resolution of $500\times 500$ pixels. The dataset used in this study is available at the following Google Drive link: https://drive.google.com/drive/folders/1JHutEYjR6tAvXNnJuMhVgJabNC8tq7z7?usp=sharing.
Algorithm 1 Pseudo-code of the Proposed Classification Model
0:
  • $\bullet$
[…]
  • $\bullet$
    $N_{\text{frames}}$: Number of frames per video
  • $\bullet$
    $S_{\text{target}}$: Target size for frame resizing
  • $\bullet$
    $p_{\text{aug}}$: Data augmentation probability
  • $\bullet$
    Batch size $=B$, Number of epochs $=E$
0:  Trained 3DCNN or ConvLSTM2D model; performance metrics
1:  1. Data Preparation
2:  for $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{test}}$ do
[…]
4:     if $(V_{i},l_{i})\in\mathcal{D}_{\text{train}}$ then
5:        $(X,y)\leftarrow(X\cup\{F_{i}\},y\cup\{l_{i}\})$
6:     else
7:        $(X_{\text{test}},y_{\text{test}})\leftarrow(X_{\text{test}}\cup\{F_{i}\},y_{\text{test}}\cup\{l_{i}\})$
8:     end if
9:  end for
10:  Apply horizontal flip augmentation to $X_{\text{train}}$ with probability $p_{\text{aug}}$
11:  Normalize $X_{\text{train}},X_{\text{test}}$ and one-hot encode labels
12:  2. Model Training
13:  $M\leftarrow\text{CreateModel}(N_{\text{frames}},S_{\text{target}})$
14:  $M\leftarrow\text{TrainModel}(M,X_{\text{train}},y_{\text{train}},B,E)$
15:  3. Model Evaluation
16:  $\hat{y}\leftarrow M(X_{\text{test}})$
17:  metrics,confusion $\text{metrics},\text{confusion matrix}\leftarrow\text{Evaluate}(y_{\text{test}},\hat{y})$
18:  return  $M$, metrics, confusion matrix
II-B Data used in this study
TABLE II: Basic information of the dataset used in this study compared with the most relevant datasets in the literature
Study Train Frames Test Frames Test Environment Focused Activity
@@ II-B Data used in this study
5 600 120 Corridor after the milking parlor Intensive farming
@@
This study 1500 500 Indoor and outdoor environment Semi-intensive farming
@@
Table II provides an overview of the data used in this research alongside datasets from previous cattle lameness detection studies. The table compares these studies based on the number of training and testing frames, the test environment, and the focused activity. As previously mentioned, the dataset consists of a total of 50 videos, including 25 normal and 25 lame cattle videos. Out of these, 30 videos (15 normal and 15 lame) were used for training the model, while the remaining 20 videos (10 normal and 10 lame) were used for testing. Previous studies focused on intensive farming environments, where cattle were primarily managed in indoor grazing conditions and data was collected from specific, limited locations such as within a barn, a passing alley, a narrow path, and corridors related to milking parlors. In contrast, this study focuses on a semi-intensive farming setting and incorporates both indoor and outdoor environments. This approach captures a wider range of cattle movement behaviors, lighting conditions, and surface variations, potentially leading to a more generalizable model for cattle lameness detection.
II-B1 Data pre-processing
@@
Preprocessing in this investigation was conducted in several distinct steps, including frame extraction and data augmentation. The overall average frame rate was approximately 35.12 frames per second. However, due to the high volume of frames, only 25 images were randomly selected from each training video, resulting in a total of 750 images. These images were then cropped to a 224×224 pixel ratio. However, data augmentation is a widely recognized method for addressing limited data issues and helps to create robust and sustainable classification models. Numerous data augmentation techniques are available, as demonstrated by Jia et al. [3], who employed nine different methods, such as Mixup, Cutout, CutMix, etc. This study considered only image flipping as an augmentation technique, applied solely to the training data. All 30 training videos (750 frames) were virtually flipped, generating a total of 1500 images. In test data, 500 images were extracted from 20 videos. However, no data augmentation was performed on the test dataset to preserve integrity and ensure a fair assessment of the model’s generalization ability.
II-C Model Construction
@@
II-C1 3D CNN
CNNs are widely used deep learning models for image processing tasks. 3D CNNs are an extension of these and can handle 3D data containing spatiotemporal information, such as videos. They utilize 3D convolutional kernels, which enable them to focus on spatial features extracted from input data and distinguish themselves from traditional 2D CNNs. Because of their ability to learn both spatial and temporal features, this study employed a 3D CNN model using the Keras Sequential API to perform binary classification, where movement patterns are assessed over time. The network accepts input tensors of shape $(25,224,224,3)$, where 25 denotes the number of consecutive frames, $224\times 224$ represents the spatial resolution of each frame, and 3 corresponds to the RGB color channels.
As the Figure 2 shows, the model architecture begins with a 3D convolutional layer comprising 32 filters with a kernel size of $(3\times 3\times 3)$, employing a ReLU activation function and same padding to preserve the original spatial and temporal dimensions. This is followed by a 3D max pooling layer with a pooling size of $(2\times 2\times 2)$ to reduce dimensionality and computational load. A second convolutional layer with 64 filters of size $(3\times 3\times 3)$ and ReLU activation is subsequently applied, again followed by a $(2\times 2\times 2)$ max pooling operation. The extracted feature maps are then flattened into a one-dimensional vector, which is passed through two fully connected layers with 128 and 64 neurons, respectively, each utilizing a ReLU activation function. To reduce the risk of overfitting, dropout layers with a dropout rate of 0.5 are inserted after each dense layer. The final output layer consists of a single neuron with a sigmoid activation function, yielding a probability score for binary classification. The model is compiled using the Adam optimizer, binary cross-entropy as the loss function, and accuracy as the evaluation metric.
Figure 2: Architectures of the 3D CNN and ConvLSTM2D models used for the cattle classification.
II-C2 ConvLSTM2D
ConvLSTM2D is a deep learning architecture that integrates the strengths of spatial feature extraction and temporal dependency modeling through the combination of CNNs and LSTM networks, respectively. This makes it particularly effective for understanding patterns and movements in sequential data. Followed by the previous model, the input layer is similar, processing video frames in the same format. In Figure 2, model 2 employed in this study begins with a ConvLSTM2D layer. Following the ConvLSTM2D layer, two successive 3D max pooling layers with pooling sizes of $(1\times 2\times 2)$ are applied to downsample the spatial dimensions while maintaining the temporal dimension. The output is then flattened into a one-dimensional feature vector. A fully connected dense layer with 128 neurons and ReLU activation is employed for high-level feature abstraction, followed by a dropout layer with a rate of 0.25 to reduce overfitting. Finally, a dense layer with sigmoid activation is used to classify cattle into ‘normal’ and ‘lame’ categories. The compilation process also follows the same procedure as in the previous model.
Figure 3: Confusion matrix of both models under the same test data.
II-D Model Evaluation
@@ III Results
The models generate predictions for each individual frame within each video in the test data. It is assumed that if the model correctly predicts the majority of frames in a video, then the entire video is considered correctly classified, despite the presence of a few incorrect frame-level predictions. The confusion matrices for both models are presented in Figure 3. As shown in Figures 3(a) and 3(b), a small number of misclassifications were observed, including 1 false positives and 1 false negative for 3D CNN, and 1 false positives and 2 false negatives for ConvLSTM2D.
The 3D CNN model demonstrated notable performance in classifying the cattle data. Table III indicates an accuracy of 90%, along with strong average precision, recall, and F1-scores. On the other hand, ConvLSTM2D achieved comparatively lower performance, with accuracy, precision, recall, and F1-score values of 85%, 90%, 81.82%, and 85.71%, respectively. Overall, the confusion matrices and classification reports highlight the models’ effectiveness in classifying healthy and defective cattle.
TABLE III: Lameness classification results of proposed 3D CNN and ConvLSTM2D
@@ IV Discussion
TABLE IV: Performance comparison of this study with existing cattle lameness classification models
@@
  • – No specific data was found
@@
Table IV presents a performance comparison between eight prior studies and the proposed model. The table evaluates models based on accuracy, precision, recall, and f1-score. The final row of the table displays the results of this study, which employed a 3D CNN model, while the preceding rows summarize the findings from other relevant studies that utilized various machine learning and deep learning classifiers. The model used in this study achieved a notable accuracy of 90%. More significantly, it demonstrated the highest precision among the studies where precision values were available, reaching 90.9%. This indicates that the model exhibits a strong ability to minimize false positives, correctly identifying positive cases with high confidence. Additionally, the 3D CNN model also yielded a robust F1-score of 90.91%. When compared to other approaches, the 3D CNN model in this study demonstrated competitive performance, whereas study [15] employed C3D-ConvLSTM achieved a slightly higher accuracy of 90.32% but a lower precision of 90.62%.
@@ V Conclusions
This paper presents a cattle lameness video dataset and deep learning-based methods for effectively distinguishing between lame and non-lame cattle from video data. Compared to existing datasets used for cattle lameness detection, this study collected data entirely from online sources, providing more comprehensive information across diverse environments, scenes, and activities. Additionally, two deep learning models, 3D CNN and ConvLSTM2D, were used to classify normal and lame cattle. From the dataset, 1,500 images were extracted for training the models, while 500 images were used for validation. The experimental results demonstrated improved or at least comparable performance to existing cattle lameness classification models. In conclusion, this study serves as a reference for detecting cattle lameness or behavior without relying on traditional multi-step approaches.